Circular No. 8231

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ja IN NGC 846 Further to IAUC 8230, M. Moore and W. Li report the LOSS/KAIT

discovery, on images taken on Oct. 22.4 and 23.4 UT, of an apparent supernova (mag ∼ 17.1) located at α = 2^{h}12^{m}11^{s}.81, δ = +44^{o}34′20′′.8 (equinox 2000.0), which is 5′′.6 west and 13′′.7 north of the nucleus of NGC 846. A KAIT image taken on Oct. 7.4 showed nothing at this position (limiting mag ∼ 19.0). A. V. Filippenko, R. Chornock, and S. Jha, University of California, Berkeley, report that inspection of CCD spectra (range 320–1000 nm), obtained on Oct. 23 with the Shane 3-m telescope at Lick Observatory, reveals that SN 2003ja is a supernova, probably of type II shortly after the explosion. The continuum is very blue and nearly featureless.

SUPERNOVAE 2003hr, 2003ix, 2003iy, AND 2003iz Filippenko, Chornock, and Jha add that inspection of CCD spectra,

obtained as above on Oct. 23 UT, reveals that SN 2003ix (IAUC 8229) is of type Ia, a few weeks past maximum brightness. SN 2003iz (cf. IAUC 8230) is also of type Ia, roughly near maximum brightness. SN 2003hr (IAUC 8191) is of type II, several months past maximum; the emission component of Hα greatly dominates over absorption, although classical P-Cyg profiles of Hβ, Na D, and Fe II are still prominent. A noisy spectrum of SN 2003iy (cf. IAUC 8230) suggests that it is a young type-II supernova, with a blue continuum and a hint of hydrogen Balmer lines.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003iz, obtained by P. Berlind on Oct. 23.28 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova after maximum. Adopting the NED recession velocity of 14453 km/s for the host galaxy, the supernova expansion velocity is 12000 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 6 ± 2 days after maximum light.

NOVAE IN M31 F. Di Mille, S. Ciroi, V. Botte, and C. S. Boschetti, Dipartimento di

Astronomia, Universitá di Padova, confirm the nova in M31 reported on IAUC 8226. Inspection of a CCD spectrum (range 380–750 nm; resolution 2.6 nm), taken at the Asiago 1.82-m telescope (+ AFOSC) on Oct. 17.948 UT, shows that hydrogen Balmer emission lines are present; the FWHM of Hα is ∼ 900 km/s. A spectrum of the Hornoch nova reported on IAUC 8222 shows strong Hα (FWHM ∼ 1100 km/s).

2003 October 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green